Using one of the laws of exponents, prove that any number raised to the 0'th power equals 1.
Answer by venugopalramana(3286)

Using one of the laws of exponents, prove that any number raised to the 0'th power equals 1.
WE KNOW X^M/X^N=(X*X*X*....M TIMES)/(X*X*X.....N TIMES)=X^(M-N)
IF WE TAKE M=N IN THIS WE GET
X^M/X^M=(X*X*X*....M TIMES)/(X*X*X*.....M TIMES)=1=X(M-M)=X^0
HENCE X^0=1